OpenStudy (anonymous):

find an equation for the tangent line to the graph of f(x)=2cos(x)-sec(x) at the point (1/3pi,f(1/3pi))

OpenStudy (baldymcgee6):

So the equation of the tangent line is simply the derivative, do you know how to take the derivative of this function?

OpenStudy (anonymous):

is it -(2+cos(2 x)) sec(x) tan(x)

OpenStudy (baldymcgee6):

i'm afraid not

OpenStudy (baldymcgee6):

what is the derivative of cos(x)

OpenStudy (anonymous):

the derivative of f(x)=2cos(x)-sec(x) is -(2+cos(2 x)) sec(x) tan(x)

OpenStudy (anonymous):

-sinx

OpenStudy (baldymcgee6):

did you actually compute the derivative or did you just get it off the internet?

OpenStudy (anonymous):

the internet

OpenStudy (anonymous):

wolframalpha is trusted

OpenStudy (baldymcgee6):

haha, that is the derivative.. But you don't know how they got that derivative.

OpenStudy (baldymcgee6):

lets do it together.

OpenStudy (anonymous):

ok...

OpenStudy (baldymcgee6):

\[ f(x)=2\cos(x)-\sec(x)\]\[ f'(x)=(2\cos(x))'-(\sec(x))'\]

OpenStudy (baldymcgee6):

that 2 is just a constant so it will stay there.. all we need to compute is the derivative of cosx and secx

OpenStudy (baldymcgee6):

\[f'(x) = 2*(-\sin(x)) - (secx*tanx)\]

OpenStudy (baldymcgee6):

\[f'(x) = 2*\sin(x) + (secx*tanx)\]

OpenStudy (anonymous):

ok..

OpenStudy (baldymcgee6):

Do you get that?

OpenStudy (anonymous):

that still does not make an equation for 1/3pi,f1/3pi

OpenStudy (baldymcgee6):

we're not done yet

OpenStudy (baldymcgee6):

first we need to find out what f(pi/3) is

OpenStudy (anonymous):

y

OpenStudy (baldymcgee6):

so we have our derivative, now to need to see what it is doing at the point (pi/3, f(pi/3))

OpenStudy (baldymcgee6):

f(x)=2cos(x)-sec(x) f(pi/3) = 2 cos(pi/3) - sec(pi/3) = 2cos(pi/3) - 1/cos(pi/3)

OpenStudy (anonymous):

ok

OpenStudy (baldymcgee6):

which = -1, so now we know our two points to evaluate our derivative at.. (pi/3, -1)

OpenStudy (anonymous):

not to be rude and i really appreciate the explaination but i need to turn in the assignment i have tutoring tomorrow so i can get the explaination then

OpenStudy (anonymous):

so do you think the answer is y-3sqrt3=1-sqrt3

OpenStudy (baldymcgee6):

the answer is -3sqrt(3)

OpenStudy (anonymous):

you have to make an equation...

OpenStudy (anonymous):

#10

OpenStudy (anonymous):

which answer choice do u believe it is

OpenStudy (baldymcgee6):

hmm.. not sure.

OpenStudy (anonymous):

wow ok